# Efficient Protein Ground State Energy Computation via Fragmentation and Reassembly

### v2

@@ VII Results
| Gly-Gly | -483.23779 | -483.25712 | 0.00400 |
| Gly-Ala | -521.82046 | -521.83696 | 0.00316 |
| Gly-Ser | -595.64593 | -595.66155 | 0.00262 |
| Carnosine (Ala-His) | -781.24420 | -781.24987 | 0.00073 |
| Aspartame (Asp-Phe) | -1010.80954 | -1011.31537 | 0.05002 |
| Cystine (Cys-Cys) | -1420.58943 | -1420.59710 | 0.00054 |
| Leu-Thr | -788.53834 | -788.55722 | 0.00232 |
| Thr-Lys | -842.85552 | -842.87107 | 0.00188 |
| Trp-His | -1137.12083 | -1137.13229 | 0.00101 |
| Phe-Ile | -902.88517 | -902.29861 | 0.06501 |
| Arg-Met | -1308.21468 | -1308.22802 | 0.00102 |
| Ser-Cys | -1027.39113 | -1027.40734 | 0.00158 |
| Tyr-Asp | -1046.06184 | -1046.07718 | 0.00147 |
| Glu-Gly | -745.47664 | -745.49541 | 0.00252 |
| His-Arg-Val | -1378.59438 | -1378.61265 | 0.00133 |
| Val-Asp-Ser | -1139.16790 | -1139.19876 | 0.00271 |
| Gly-His-Lys | -1155.42028 | -1155.44473 | 0.00212 |
| Val-Ala-Ser | -954.09354 | -954.12968 | 0.00386 |
| Gly-Val-Ala | -841.68835 | -841.71855 | 0.00358 |
| Ser-Gly-Glu | -1062.00937 | -1062.04545 | 0.00340 |
Table 2 presents the results accounting for the fragmentation of the amino acids into their corresponding four different groups presented in Figure 3. In this case, the error is higher due to the further fragmentation of the molecules but still very promising (mean %RE of ), with only considerable differences in the Asparagine amino acid (%RE of ).
@@
| Histidine | -538.52442 | -537.58932 | 0.17394 |
| Leucine | -433.42225 | -434.01054 | 0.13555 |
| Isoleucine | -432.82708 | -434.01055 | 0.27268 |
| Lysine | -487.74060 | -487.36827 | 0.07639 |
| Methionine | -788.02138 | -787.09064 | 0.11825 |
| Phenylalanine | -544.43743 | -544.04328 | 0.07245 |
| Threonine | -430.09637 | -429.12416 | 0.22656 |
| Tryptophan | -673.57377 | -673.15017 | 0.06293 |
| Valine | -394.84749 | -394.45688 | 0.09902 |
| Arginine | -595.17254 | -594.18510 | 0.16618 |
| Cysteine | -710.85730 | -715.26885 | 0.61677 |
| Glutamine | -521.82178 | -516.80354 | 0.97101 |
| Asparagine | -519.42952 | -479.61163 | 8.30211 |
| Tyrosine | -618.27595 | -611.07066 | 1.17913 |
| Serine | -391.51594 | -391.12359 | 0.10031 |
| Glycine | -279.11151 | -278.67407 | 0.15697 |
| Aspartic acid | -502.76713 | -502.31178 | 0.09065 |
| Glutamic acid | -541.34980 | -540.89235 | 0.08457 |
| Proline | -393.70020 | -393.87365 | 0.04436 |
| Alanine | -317.69135 | -317.28419 | 0.12833 |
@@ VIII Discussion
The presented approach involved summing the energy of each amino acid and adding corrections by incorporating or subtracting molecules involved in the formation of the original peptide. Remarkably, this method proved to be effective even in cases involving uncommon bond formations. For example, cystine, a dimer formed by two cysteines, is typically bonded through a sulfur link instead of the carboxyl and amino terminal groups. In such cases, we subtracted the GSE of two hydrogen molecules () rather than a water molecule during the reassembly process, still obtaining highly accurate results. Similarly, in the case of aspartame, a dimer consisting of asparagine and phenylalanine, a methyl group () is added during the formation of the molecule. To account for this, we computed the GSE of and added it to the final energy, again achieving very high accuracy. This is of utmost importance since, unlike previous methodologies, we provide a very easy pipeline to implement [16].
Similarly, the accuracy was almost maintained even for amino acid fragmentation, where molecules were completely divided into much smaller fragments. Despite this, our methodology appears to effectively reconstitute the properties of the original molecule without significant losses. Interestingly, this was true for almost all the side chains of the amino acids, even though some were much more complex than others. However, asparagine exhibited a higher error rate, indicating that our approach might be missing some interactions in this specific case.

### v3

@@ VII Results
| Gly-Gly | -483.23779 | -483.25713 | 0.00400 |
| Gly-Ala | -521.82046 | -521.83697 | 0.00317 |
| Gly-Ser | -595.64593 | -595.66156 | 0.00262 |
| Carnosine (Ala-His) | -781.24420 | -781.25935 | 0.00194 |
| Aspartame (Asp-Phe) | -1010.80954 | -1011.31538 | 0.05004 |
| Cystine (Cys-Cys) | -1420.58943 | -1420.59711 | 0.00054 |
| Leu-Thr | -788.53834 | -788.55273 | 0.00182 |
| Thr-Lys | -842.85520 | -842.87108 | 0.00188 |
| Trp-His | -1137.12083 | -1137.14177 | 0.00184 |
| Phe-Ile | -902.88517 | -902.89959 | 0.00160 |
| Arg-Met | -1308.21468 | -1308.22803 | 0.00102 |
| Ser-Cys | -1027.39113 | -1027.40735 | 0.00158 |
| Tyr-Asp | -1046.06184 | -1046.07719 | 0.00147 |
| Glu-Gly | -745.47664 | -745.49542 | 0.00252 |
| His-Arg-Val | -1378.59438 | -1378.62213 | 0.00201 |
| Val-Asp-Ser | -1139.16790 | -1139.19877 | 0.00271 |
| Gly-His-Lys | -1155.42028 | -1155.45421 | 0.00294 |
| Val-Ala-Ser | -954.09354 | -954.12300 | 0.00309 |
| Gly-Val-Ala | -841.68835 | -841.71857 | 0.00359 |
| Ser-Gly-Glu | -1062.00937 | -1062.04546 | 0.00340 |
Table 2 presents the results accounting for the fragmentation of the amino acids into their corresponding four different groups presented in Figure 3. In this case, the error is higher due to the further fragmentation of the molecules but still very promising (mean %RE of ).
@@
| Histidine | -538.53389 | -537.58932 | 0.17540 |
| Leucine | -433.42225 | -434.01055 | 0.13573 |
| Isoleucine | -433.42805 | -434.01055 | 0.13439 |
| Lysine | -487.74061 | -487.36827 | 0.076339 |
| Methionine | -788.02139 | -787.09064 | 0.11811 |
| Phenylalanine | -544.43743 | -544.04328 | 0.072395 |
| Threonine | -430.09637 | -429.12416 | 0.22604 |
| Tryptophan | -673.57378 | -673.15017 | 0.062891 |
| Valine | -394.84750 | -394.45688 | 0.098928 |
| Arginine | -595.17255 | -594.18510 | 0.16591 |
| Cysteine | -710.85730 | -715.26885 | 0.62060 |
| Glutamine | -521.82179 | -516.80354 | 0.96168 |
| Asparagine | -483.23923 | -479.61163 | 0.75068 |
| Tyrosine | -618.27595 | -611.07066 | 1.1654 |
| Serine | -391.51594 | -391.12360 | 0.10021 |
| Glycine | -279.11151 | -278.67407 | 0.15673 |
| Aspartic acid | -502.76713 | -502.31178 | 0.090569 |
| Glutamic acid | -541.34980 | -540.89235 | 0.084502 |
| Proline | -393.70020 | -393.87365 | 0.044055 |
| Alanine | -317.69136 | -317.28420 | 0.12816 |
@@ VIII Discussion
The presented approach involved summing the energy of each amino acid and adding corrections by incorporating or subtracting molecules involved in the formation of the original peptide. Remarkably, this method proved to be effective even in cases involving uncommon bond formations. For example, cystine, a dimer formed by two cysteines, is typically bonded through a sulfur link instead of the carboxyl and amino terminal groups. In such cases, we subtracted the GSE of two hydrogen molecules () rather than a water molecule during the reassembly process, still obtaining highly accurate results. Similarly, in the case of aspartame, a dimer consisting of aspartic acid and phenylalanine, a methyl group () is added during the formation of the molecule. To account for this, we computed the GSE of and added it to the final energy, again achieving very high accuracy. This is of utmost importance since, unlike previous methodologies, we provide a very easy pipeline to implement [16].
Similarly, the accuracy was almost maintained even for amino acid fragmentation, where molecules were completely divided into much smaller fragments. Despite this, our methodology appears to effectively reconstitute the properties of the original molecule without significant losses. Interestingly, this was true for almost all side chains of amino acids, even though some were much more complex than others.